yane is 3 years older than tim and twice as old as henry.tim is 2 years older than henry. how old is each.

Answer by Timnewman(323)

hi dear,
Let Y=yane,
H=henry,
T=Tim
From your first sentence,
y=t+6----(1)
y=2h---(2)
From your second sentence,
t=h+2----(3)
Equate equation 1 and 2
t+6=2h----(4)
put Equ 3 in 4
h+2+6=2h
h+8=2h
2h-h=8
h=8
put h=8 in3
t=8+2
t=10
put t=10 in equ 1
y=10+6
Y=16
The ages of yane,tim,henry is 16yrs,10yrs and 8yrs respectively.
HOPE THIS HELPS?
::TIMNEWMAN::